# Optimal energy storage in the Tavis-Cummings quantum battery

### v1

@@ IV Discussion
In this work, we have investigated the realization of optimal energy storage in the TC battery. By selecting the Fock state as the initial state of the charger, we have introduced the invariant subspace method to investigate the quantum dynamics of the TC model. This approach has enabled us to demonstrate that SU(2) symmetry emerges in case where either $n_{0}\gg N_{b}$ or $N_{b}\gg n_{0}$. Consequently, under these two conditions, the optimal energy storage in the TC battery can be achieved. In the scenario where $N_{b}\!=\!n_{0}$, we have conducted numerical investigations into the energy storage problem, confirming the existence of a negative relation between the battery-charger entanglement and the energy storage capacity. Our numerical results have indicated the possibility of asymptotically achieving optimal energy storage in the limit $N_{b}\!=\!n_{0}\!\gg\!1$.
It is worth noting that the TC battery is possibly realized using the state-of-the-art solid-state technology through the circuit/cavity QED Fink et al. (2009); Yang et al. (2020); Leek et al. (2009); Lolli et al. (2015); Liu et al. (2023). Reference Fink et al. (2009) reported an ideal realization of the TC model in the absence of atom number and coupling fluctuations by embedding a discrete number of fully controllable superconducting qubits at fixed positions into a transmission line resonator. Reference Liu et al. (2023) realized strong coupling between deterministic single-atom arrays and a high-finesse miniature optical cavity. While our theoretical starting point has been the Fock state as the initial charge, in experimental settings, coherent light can be used as an approximation to the Fock state. This simplifies the difficulty of experimentally preparing Fock states. Consequently, our theoretical approaches to the optimal energy storage can be implemented in a well-designed circuit/cavity QED system. Our work not only contributes to a profound understanding of the algebraic structure of the TC model but also offers significant insights into the problem of the optimal energy storage. We anticipate that our approach can be extended to address other quantum battery-related challenges in the future.
@@ Acknowledgements.
This work was supported by the NSFC (Grants No. 12275215, No. 12305028, No. 12247103, and No. 12175178), the Major Basic Research Program of Natural Science of Shaanxi Province (Grant No. 2021JCW-19), Shaanxi Fundamental Science Research Project for Mathematics and Physics (Grant No. 22JSZ005) and the Youth Innovation Team of Shaanxi Universities. Q.-K. Wan and H.-L. Shi received partial support from the NSFC key grants No. 12134015 and No. 92365202. H.-L. Shi was supported by the European Commission through the H2020 QuantERA ERA-NET Cofund in Quantum Technologies project “MENTA”. The authors H.-L. Shi and H.-Y. Yang contributed equally to this work.
@@
According to the algebraic relations:
@@
The spin-$d$ irreducible representation of the generators of su(2) algebra is given by
$\bm{S}^{x}=\frac{1}{2}\begin{pmatrix}0&f_{1}&\\ f_{1}&0&f_{2}&\\ &\ddots&\ddots&\ddots&\\ &&f_{d-1}&0&f_{d}\\ &&&f_{d}&0\end{pmatrix},$ (44)
$\bm{S}^{y}=\frac{1}{2\rm i}\begin{pmatrix}0&-f_{1}&\\ f_{1}&0&-f_{2}&\\ &\ddots&\ddots&\ddots&\\ &&f_{d-1}&0&-f_{d}\\ &&&f_{d}&0\end{pmatrix},$ (50)
@@
where $f_{j}\equiv\sqrt{j(2d-j+1)}$. If we take $d\!=\!N_{b}/2$, then $\bm{S}^{\alpha}$ ($\alpha\!=\!x,y,z$) correspond to the operators $\bm{J}^{\alpha}$, see Eqs. (28), (30), and (31). If we take $d\!=\!n_{0}/2$, then $\bm{S}^{\alpha}$ correspond to the operators $\bm{\tilde{J}}^{\alpha}$, see Eqs. (33) and (35).

### v3

@@ IV Discussion
In this paper, we have investigated the realization of optimal energy storage in the TC battery. By selecting the Fock state as the initial state of the charger, we have introduced the invariant subspace method to investigate the quantum dynamics of the TC model. This approach has enabled us to demonstrate that SU(2) symmetry emerges in cases where either $n_{0}\gg N_{b}$ or $N_{b}\gg n_{0}$. Consequently, under these two conditions, optimal energy storage in the TC battery can be achieved. In the scenario where $N_{b}\!=\!n_{0}$, we have conducted numerical investigations into the energy storage problem, confirming the existence of a negative relation between the battery-charger entanglement and the energy storage capacity. Our numerical results have indicated the possibility of asymptotically achieving optimal energy storage in the limit $N_{b}\!=\!n_{0}\!\gg\!1$.
It is worth noting that the TC battery is possibly realized using the state-of-the-art solid-state technology through the circuit/cavity QED Fink et al. (2009); Yang et al. (2020); Leek et al. (2009); Lolli et al. (2015); Liu et al. (2023). Reference Fink et al. (2009) reported an ideal realization of the TC model in the absence of atom number and coupling fluctuations by embedding a discrete number of fully controllable superconducting qubits at fixed positions into a transmission line resonator. Reference Liu et al. (2023) realized strong coupling between deterministic single-atom arrays and a high-finesse miniature optical cavity. While our theoretical starting point has been the Fock state as the initial charge, in experimental settings, coherent light can be used as an approximation to the Fock state. This simplifies the difficulty of experimentally preparing Fock states. Consequently, our theoretical approaches to optimal energy storage can be implemented in a well-designed circuit/cavity QED system. Our work not only contributes to a profound understanding of the algebraic structure of the TC model but also offers significant insights into the problem of optimal energy storage. We anticipate that our approach can be extended to address other quantum battery-related challenges in the future.
@@ Acknowledgements.
This work was supported by the NSFC (Grants No. 12275215, No. 12305028, No. 12247103, and No. 12175178), the Major Basic Research Program of Natural Science of Shaanxi Province (Grant No. 2021JCW-19), Shaanxi Fundamental Science Research Project for Mathematics and Physics (Grant No. 22JSZ005) and the Youth Innovation Team of Shaanxi Universities. Q.-K.W. and H.-L.S. received partial support from the NSFC Key Grants No. 12134015 and No. 92365202. H.-L.S. was supported by the European Commission through the H2020 QuantERA ERA-NET Cofund in Quantum Technologies project “MENTA.” H.-L.S. and H.-Y.Y. contributed equally to this work.
@@
According to the algebraic relations,
@@
[…]
$\bm{S}^{x}=\frac{1}{2}\begin{pmatrix}0&f_{1}&\\ f_{1}&0&f_{2}&\\ &\ddots&\ddots&\ddots&\\ &&f_{d-1}&0&f_{d}\\ &&&f_{d}&0\end{pmatrix},$
$\bm{S}^{y}=\frac{1}{2\rm i}\begin{pmatrix}0&-f_{1}&\\ f_{1}&0&-f_{2}&\\ &\ddots&\ddots&\ddots&\\ &&f_{d-1}&0&-f_{d}\\ &&&f_{d}&0\end{pmatrix},$
@@
where $f_{j}\equiv\sqrt{j(2d-j+1)}$. If we take $d\!=\!N_{b}/2$, then $\bm{S}^{\alpha}$ ($\alpha\!=\!x,y,z$) correspond to the operators $\bm{J}^{\alpha}$ [see Eqs.  (16),  (18), and (19)]. If we take $d\!=\!n_{0}/2$, then $\bm{S}^{\alpha}$ correspond to the operators $\bm{\tilde{J}}^{\alpha}$ [see Eqs.  (21) and (23)].